OpenStudy (zyberg):

Could anyone explain to me how to convert number to their negative bases? (There is one thing that I don't understand about calculation algorithm in the comments)

OpenStudy (zyberg):

Negative bases wiki shows us an example of how t calculate 146 in negaternary: /(146/-3 = -48 remainder 2/) /(-48/-3 = 16 remainder 0/) /(16/-3 =-5remainder1/) /(-5/-3 =2 remainder1/)What is this? How is it possible, why is that not a mistake? /(2/-3 = 0remainder2/) I had encountered this same "division error" in a few places, so perhaps it is intentional? @ganeshie8 maybe you know anything about this?

ganeshie8 (ganeshie8):